Mary has $2.20 in dimes and nickels. If she has 10 more dimes than nickels, how many of each coin does she have?
I've set this up a couple different ways with no luck...Thanks!

Answer by Edwin McCravy(20062)

Mary has $2.20 in dimes and nickels. If she has 10 more dimes than nickels, how many of each coin does she have?
coins      number of     value
--------------------------------
dimes         n+10     .10(n+10) 
nickels        n          .05n
--------------------------------
total                     2.20

.10(n+10) + .05n = 2.20

Multiply through by 100

   10(n+10) + 5n = 220  
  10n + 100 + 5n = 220                  
       15n + 100 = 220
             15n = 120
               n = 8

She has 8 nickels.

she has 10 more dimes than nickels,

So she has 8+10 or 18 dimes.

Checking:
8 nickels is 40 cents and 18 dimes is $1.80,
and 40 cents plus $1.80 is $2.20.

Edwin